OpenStudy (anonymous):

1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4

OpenStudy (helder_edwin):

x=4 is a vertical so it has no slope (or it's equal to infinty) so the line through (5,-2) has to be vertical as well

OpenStudy (anonymous):

So does that mean there is no line that can be parallel to (5,-2) @helder_edwin

OpenStudy (anonymous):

parallel line have same gradient ,the line x=4 has a slope of zero which means the line which passes through(5,-2) has slope 0,since y=mx+c,then -2=0*5+c,which implies c=-2,then the equation of the line is x=-2

OpenStudy (helder_edwin):

no! vertical lines have the form \[ x=z \] where z is any number @juan_chi is right with the answer though x=4 DOES NOT have slope zero

OpenStudy (helder_edwin):

sorry the answer x=-2 is not right

OpenStudy (helder_edwin):

plus remember the ordered pair is (x,y)

OpenStudy (anonymous):

So what would be the answer ?

OpenStudy (helder_edwin):

come on \[ (x,y)=(5,-2) \]

OpenStudy (anonymous):

So would the answer be 5=-2x+4 ?

OpenStudy (helder_edwin):

no \[ (a,b)=(c,d)\quad \text{if and only if}\quad a=c\wedge b=d \]

OpenStudy (anonymous):

Ok so x=5

OpenStudy (helder_edwin):

yes!

OpenStudy (anonymous):

Thank You !

OpenStudy (helder_edwin):

sure